OpenStudy (anonymous):

whats the following arithmetic sequences -4,2,0,-2 1)tn=4+2(n-1) 2)tn=2+4(n-1) 3)tn=1+4(n-2) 4)tn=4-2(n-1)

OpenStudy (anonymous):

the arithmetic sequence is ..... 4,2,0,-2..... first term is 4 common difference is 2-4=-2 a=4 d=-2 in an arithmetic sequence \[t_n=a+(n-1)d\] =\[4+(n-1)(-2)\] =4-2(n-1)